OpenStudy (anonymous):

Solve the following Initial Value Theorem: dy/dt = sec^2(t) - sin(t), y(pi/4) = 1

OpenStudy (anonymous):

The solution to the equation is: \[y(t) = c + \cos(t) + \tan(t) \] Replace each t with Pi/4. The function value y(Pi/4) is one. \[1=c+\cos{\pi \over 4} +\tan{\pi \over 4}\] Solve for c and replace it's value in the original solution. \[y(t)= -\frac{1}{\sqrt{2}}+\cos(t)+\tan(t) \]